# An Information Theoretic Perspective on Conformal Prediction

### v1

@@ F.2 Federated Setting
We run SCP with the final global model assuming calibration and test data sets at the server, or equivalently that the clients share their scores with the server. This reflects the best inefficiency results we can hope for with the global model, as in practice we might need to resort to privacy preserving methods that are likely to hurt performance. Notably, we considered the quantile of quantiles approach of Humbert et al. (2023) as well as the simpler alternative proposed in Lu & Kalpathy-Cramer (2021), but in both cases we got varying degrees of coverage due to the data heterogeneity among devices introduced in the distribution-based label imbalance setup. Addressing these shortcomings is a promising avenue for future research for conformal prediction in the federated setting.
@@
After learning the global model, we fine tune it on the local training data of each device to obtain a personalized model. We then run SCP individually for each device with local calibration and test data sets and report the average prediction sets across all clients. Importantly, since each client has access to only a small number of data points, we do not always achieve valid coverage in the personalized setting. More precisely, all personalized models on CIFAR10 and EMNIST achieved marginal coverage of around 97%, while for CIFAR100 that value dropped to 90%. Nonetheless, all methods get similar coverages so the results remain comparable. Interestingly, even with personalized models, which implicitly already estimate $Q_{Y|X,Z}$, updating the personalized model as in (8) with the global head $Q_{Z|X,Y}$, still results in non-negligible improvements in performance, as shown in Tables 11 and 12, as well as 13 and 14.
Table 11: Inefficiency of ConfTr Prediction with thresholding (THR) for CIFAR10 with different levels of data heterogeneity among clients both global (GLO) and personalized (PER) models in the federated setting for a target $\alpha=0.01$.
@@
GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$
@@
Table 12: Inefficiency of ConfTr Prediction with APS for CIFAR10 with different levels of data heterogeneity among clients both global (GLO) and personalized (PER) models in the federated setting for a target $\alpha=0.01$.
@@
GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$
@@
Table 13: Inefficiency of Conformal Prediction with thresholding (THR) for each dataset with both global (GLO) and personalized (PER) models in the federated setting for a target $\alpha=0.01$.
@@
GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$
@@
ConfTr + class $2.88_{\pm 0.02}$ $\mathbf{1.98_{\pm 0.02}}$ $\mathbf{1.69_{\pm 0.31}}$ $\mathbf{1.55_{\pm 0.24}}$ $3.53_{\pm 0.09}$ $3.39_{\pm 0.08}$ $2.86_{\pm 0.38}$ $2.79_{\pm 0.35}$ $58.53_{\pm 1.40}$ $56.03_{\pm 1.29}$ $\mathbf{25.22_{\pm 7.40}}$ $\mathbf{23.87_{\pm 7.02}}$
@@
Table 14: Inefficiency of Conformal Prediction with APS for each dataset with both global (GLO) and personalized (PER) models in the federated setting for a target $\alpha=0.01$.
@@
GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$ GLO GLO${}_{+\text{side}}$ PER PER${}_{+\text{side}}$
@@
F.3 Evaluating the lower bounds on the set size
In this section we will evaluate our two lower bounds on the expected $[\log|C(X)|]^{+}$. The first one can be obtained by rearranging the simple Fano bound (c.f. Proposition B.7) whereas the second one can be obtained from the model-based Fano entropy upper bounds (c.f. Theorem D.2). The main challenge in evaluating these bounds is in that we require the ground truth entropy $H(Y|X)$ or $H(Y|\hat{Y})$ which in general are not available. To proceed, we adopt the versions of the bounds that depend on $H(Y|\hat{Y})$. When $\hat{Y}$ is discrete, we can get a tractable lower bound to $H(Y|\hat{Y})$ via a maximum likelihood estimate of the entropy Paninski (2003). More specifically, we have that
@@ F.3 Evaluating the lower bounds on the set size
Based on this, we evaluate our set size bounds on ResNet models trained with CE on CIFAR 10 and CIFAR 100. As the logits $\hat{Y}$ used for CP are not discrete, we perform K-means clustering on them and construct a vector quantized $\hat{Y}_{vq}$ by assigning to each logit vector its closest cluster centroid. We then use $\hat{Y}_{vq}$ to perform CP with thresholding and also use $\hat{Y}_{vq}$ to obtain a lower bound on $H(Y|\hat{Y}_{vq})$ via a maximum likelihood estimate for $H(Y,\hat{Y}_{vq})$. For this maximum likelihood estimate we also perform the Miller-Madow bias correction Paninski (2003).

### v2

@@ G.2 Federated Setting
We run SCP with the final global model assuming calibration and test data sets at the server, or equivalently that the clients share their scores with the server. This reflects the best inefficiency results we can hope for with the global model, as in practice we might need to resort to privacy-preserving methods that are likely to hurt performance. Notably, we considered the quantile of quantiles approach of Humbert et al. [22] as well as the simpler alternative proposed in [35], but in both cases we got varying degrees of coverage due to the data heterogeneity among devices introduced in the distribution-based label imbalance setup. Addressing these shortcomings is a promising avenue for future research for conformal prediction in the federated setting.
@@
After learning the global model, we fine-tune it on the local training data of each device to obtain a personalized model. We then run SCP individually for each device with local calibration and test data sets and report the average prediction sets across all clients. Importantly, since each client has access to only a small number of data points, we do not always achieve valid coverage in the personalized setting. More precisely, all personalized models on CIFAR10 and EMNIST achieved marginal coverage of around 97%, while for CIFAR100 that value dropped to 90%. Nonetheless, all methods get similar coverages so the results remain comparable. Interestingly, even with personalized models, which implicitly already estimate $Q_{Y|X,Z}$, updating the personalized model as in (9) with the global head $Q_{Z|X,Y}$, still results in non-negligible improvements in performance, as shown in Tables 15 and 16, as well as 17 and 18.
Table 15: Inefficiency results for different degrees of data heterogeneity in FL with THR. Average prediction set size with THR for CIFAR10 with different levels of data heterogeneity among clients for both global (GLO) and personalized (PER) models in the federated setting for a target $\alpha=0.01$. We use ${}_{+\text{SI}}$ to indicate the inclusion of side information. Lower is better.
@@
GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$
@@
Table 16: Inefficiency results for different degrees of data heterogeneity in FL with APS. Average prediction set size with APS for CIFAR10 with different levels of data heterogeneity among clients for both global (GLO) and personalized (PER) models in the federated setting for a target $\alpha=0.01$. We use ${}_{+\text{SI}}$ to indicate the inclusion of side information. Lower is better.
@@
GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$
@@
Table 17: Inefficiency results for global (GLO) and personalized (PER) models with THR. Average prediction set size with THR for federated conformal training with a target $\alpha=0.01$. We use ${}_{+\text{SI}}$ to indicate the inclusion of side information. We show in bold all values within one standard deviation of the best result. Lower is better.
@@
GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$
@@
[…]
@@
Table 18: Inefficiency results for global (GLO) and personalized (PER) models with APS. Average prediction set size with APS for federated conformal training with a target $\alpha=0.01$. We use ${}_{+\text{SI}}$ to indicate the inclusion of side information. We show in bold all values within one standard deviation of the best result. Lower is better.
@@
GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$ GLO GLO${}_{+\text{SI}}$ PER PER${}_{+\text{SI}}$
@@
G.3 Evaluating the lower bounds on the set size
In this section we will evaluate our two lower bounds on the expected $[\log|C(X)|]^{+}$. The first one can be obtained by rearranging the simple Fano bound (c.f. Proposition C.7) whereas the second one can be obtained from the model-based Fano entropy upper bounds (c.f. Theorem E.2). The main challenge in evaluating these bounds is in that we require the ground truth entropy $H(Y|X)$ or $H(Y|\hat{Y})$ which in general are not available. To proceed, we adopt the versions of the bounds that depend on $H(Y|\hat{Y})$. When $\hat{Y}$ is discrete, we can get a tractable lower bound to $H(Y|\hat{Y})$ via a maximum likelihood estimate of the entropy [44]. More specifically, we have that
@@ G.3 Evaluating the lower bounds on the set size
Based on this, we evaluate our set size bounds on ResNet models trained with CE on CIFAR 10 and CIFAR 100. As the logits $\hat{Y}$ used for CP are not discrete, we perform K-means clustering on them and construct a vector quantized $\hat{Y}_{vq}$ by assigning to each logit vector its closest cluster centroid. We then use $\hat{Y}_{vq}$ to perform CP with thresholding and also use $\hat{Y}_{vq}$ to obtain a lower bound on $H(Y|\hat{Y}_{vq})$ via a maximum likelihood estimate for $H(Y,\hat{Y}_{vq})$. For this maximum likelihood estimate we also perform the Miller-Madow bias correction [44].